what is the square root of 2-i?
1
Expert's answer
2021-10-26T08:56:17-0400

The polar form of "2-i" is "\sqrt{5}\bigg(\cos\big(-\tan^{-1}(0.5)\big)+i\sin\big(-\tan^{-1}(0.5)\big)\bigg)".

According to the De Moivre's Formula all square roots of complex number "2-i" are given by


"\sqrt{\sqrt{5}}\bigg(\cos\big(\dfrac{-\tan^{-1}(0.5)+2\pi k}{2}\big)"

"+i\sin\big(\dfrac{-\tan^{-1}(0.5)+2\pi k}{2}\big)\bigg), k=0, 1"

"k=0:"


"\sqrt{\sqrt{5}}\bigg(\cos\big(\dfrac{-\tan^{-1}(0.5)+2\pi (0)}{2}\big)"

"+i\sin\big(\dfrac{-\tan^{-1}(0.5)+2\pi (0)}{2}\big)\bigg)""=\sqrt[4]{5}\bigg(\cos\big(\dfrac{\tan^{-1}(0.5)}{2}\big)-i\sin\big(\dfrac{\tan^{-1}(0.5)}{2}\big)\bigg)"


"k=1:"


"\sqrt{\sqrt{5}}\bigg(\cos\big(\dfrac{-\tan^{-1}(0.5)+2\pi (1)}{2}\big)"

"+i\sin\big(\dfrac{-\tan^{-1}(0.5)+2\pi (1)}{2}\big)\bigg)""=\sqrt[4]{5}\bigg(-\cos\big(\dfrac{\tan^{-1}(0.5)}{2}\big)+\sin\big(\dfrac{\tan^{-1}(0.5)}{2}\big)\bigg)"